### v1

@@ 4.1 Model
Voluntary worker flows from $j$ to $k$ are given by the measure of workers employed at $j$ ($L_{j}$), who are neither forced to reallocate (with probability $\rho_{j}$) nor forced to move to non-employment (with probability $\delta_{j}$), search for an alternative employer (with probability $\lambda_{1}$), and get an offer from employer $k$ with probability $f_{k}$.272727 The probabilities $\rho_{j}$ and $\delta_{j}$ are assumed to be mutually exclusive. In Equation (2), workers choose whether to remain with employer $j$ or move to $k$, which occurs with probability $\text{Pr}(k\succ j)$. This is the sense in which these flows are voluntary and reveal workers’ preferences over the set of employers $\mathcal{J}$. We follow Sorkin (2018) and assume there are i.i.d. taste shocks $\{\varepsilon_{i}\}_{i\in\{\mathcal{J},N\}}$ that make the value of working at employer $j$ and of non-employment worker-specific in each period. We further assume that these shocks are drawn from a Gumbel distribution with location parameter normalized to zero and scale parameter $\sigma^{-1}$. [^28] Normalizing the location parameter to zero is without loss of generality because it shifts the value of working at each employer $j$ and of being in non-employment by the same amount. With this assumption, the acceptance probability for a worker with a job offer from employer $k$ currently working at $j$ is given by
@@
We depart from Sorkin (2018) by making an explicit functional form assumption on the utility flow $u_{j}$ of working at employer $j$ [^29] Sorkin (2018) makes a similar assumption on the value $V_{j}$ of employers; Lagos (2024) also follows a similar approach. By contrast, $V_{j}$ is not additively separable into a pay and non-pay in our formulation. We assume that $u_{j}$ depends linearly on a pay component $\ln W_{j}$ and a non-pay (or amenity) component $a_{j}$
@@
The additive formulation $\ln W_{j}+a_{j}$ is standard in the literature that estimates static models of worker preferences over employers (Card et al., 2018; Lamadon et al., 2022). With an additional assumption to pin down the scale of the i.i.d taste shocks $\varepsilon_{i}\}_{i\in\{\mathcal{J},N\}}$, we can recover the non-pay component $a_{j}$ associated with working at employer $j$ by inverting the value function in (3), conditional on the wage $W_{j}$ paid by employer $j$. We discuss this assumption below.
@@
We complete the description of the model by noting that there are similar expressions for workers flows to and from non-employment (counterparts to Equation 1 and Equation 2) and for the value of worker in non-employment (a counterpart to Equation 3), where workers in non-employment get a draw from the job offer distribution $\{f_{j}\}$ at rate $\lambda_{0}$. These expressions follow naturally and are relegated to Appendix C.
@@ 4.2 Identification
The employer-level parameters $\{\sigma V_{j},L_{j},f_{j},\rho_{j},\delta_{j}\}_{j\in\mathcal{J}}$ and the transition parameters $\lambda_{0}$ and $\lambda_{1}$ are identified from (i) the realized mobility flows between employers $M_{jk}$, and (ii) which employers are shrinking or growing (see Sorkin, 2018, Section V.B). The notation $\sigma V_{j}$ emphasizes that worker flows pin down the value of working at employer $j$ up to the scale of the idiosyncratic taste shocks, which is identified in the next steps.
We briefly outline the main source of identification for each parameter intuitively and refer to the original paper for details. The main source of identification for $\sigma V_{j}$ is the aggregation of the realized mobility flows $M_{jk}$. An estimate for $L_{j}$ is obtained directly from the data as the share of worker-year observations at employer $j$. $f_{j}$ is identified from the share of non-employment flows to employer $j$. The separation shocks $\delta_{j}$ and $\rho_{j}$ are identified from the separation flows at shrinking firms, which give information on involuntary worker flows. Finally, the transition parameters $\lambda_{0}$ and $\lambda_{1}$ are pinned down by accounting for the aggregate transition flows, respectively from non-employment to employment and between employers.
Two sample restrictions are required for identification at this step. First, the aggregation of mobility flows only identifies $V_{j}$ within the strongly connected sets of employers.303030 Within a strongly connected set of employers, each employer has at least one worker moving in and one worker moving out. This condition is required for the fixed-point associated with the appropriately scaled matrix of worker flows to exist. See, e.g., Jackson (2010) for details on social network definitions. Second, each employer must hire at least one worker from non-employment for its sampling weight $f_{j}$ to be well-defined.
@@
where $X_{it}$ are exogenous worker characteristics (such as age), $\alpha_{i}$ is the unobserved worker effect, $j(i,t)$ is worker $i$’s employer in period $t$, and $e_{it}$ is an idiosyncratic shock. $\psi_{j(i,t)}$ is only identified within the connected sets of employers, which is already required to identify $V_{j}$ from worker flows. [^32] A strongly connected set is connected by definition.
Specifically, we make the assumption that employers post piece-rate wage contracts as part of the utility bundle $V_{j}$. Piece-rate contracts are a commonly used assumption in random search models with wage-posting. [^33] See, for instance, Barlevy (2008) and Engbom and Moser (2022). This assumption implies that the wage premium paid by employers is log-additive in workers’ experience ($X_{it}$), ability ($\alpha_{i}$), and a time-varying and non-persistent productivity shock ($e_{it}$). Equation (6) is then correctly specified as long as individual-specific productivity does not interact with mobility decisions. We therefore also require that (i) individual-specific productivity is log-linear in the flow value of non-employment, and (ii) time-varying productivity shocks $e_{it}$ are mean-independent conditional on the realization of i.i.d. taste shocks. These assumptions ensure that individual-specific productivity does not affect the choice to move between employers and between employment and non-employment, in line with the model of mobility described so far.
@@
Conditional on $u_{j}$ and a measure for $\ln W_{j}$, the non-pay parameters can be recovered by inverting Equation (4), which requires to pin down the parameter $\sigma$ controlling the scale of the i.i.d. taste shocks to recover $u_{j}$. We do this by assuming that the moment condition $\text{Var}(u_{j})=\text{Var}\big{(}\ln W_{j}\big{)}$ holds, so the variance in flow utility is the same as the variance in the flow utility of pay $\ln W_{j}$. An intuitive justification for this moment condition is that the flow utility of pay can be expected to be an important component of the flow utility of a job, so their variance should be of similar magnitude. However, this moment condition is not without loss of generality because it restricts the correlation between $\ln W_{j}$ and $a_{j}$ to be non-positive. It is immediate to check that $\text{Var}(u_{j})=\text{Var}\big{(}\ln W_{j}\big{)}$ implies $\text{Corr}(\ln W_{j},a_{j})\leq 0$. In this specific sense, the non-pay parameters we recover can be interpreted as having a compensating differentials component by definition.
@@ 4.3 Estimation
Our sample restrictions follow from the model structure and the conditions required for identification. We restrict the sample to the period 2015-2019 and retain workers aged 20 to 60 (inclusive). We also restrict attention to relatively larger and more stable employers. Specifically, we require each establishment to be observed in at least two separate years and to have at least five non-singleton workers on average in each year, where a non-singleton worker is a worker observed at some later point within the sample period. In addition, the identification conditions outlined in Section 4.2 impose extra restrictions on the set of establishments included in the estimation sample. We restrict the sample to the largest strongly connected set of establishments and impose that each of these establishments hire at least one individual from non-employment. [^35] Focusing on the largest connected set is common in the literature. See, among others, Card et al. (2013) and Bonhomme et al. (2020) in the context of the two-sided unobserved heterogeneity regression model (6) and Sorkin (2018) and Morchio and Moser (2024) in the context of the fixed-point revealed preferences model. These last two restrictions are interdependent, and they are imposed recursively until the set of establishments converges.
@@
We obtain estimates of the establishment wage premium from the two-sided unobserved heterogeneity regression (6). We estimate this equation separately in each sample. Our wage measure is log average earnings per hour, where earnings are total pre-tax earnings and hours are total contract hours over the duration of each yearly employment spell. The vector of worker-level characteristics $X_{it}^{\prime}$ includes year-by-education fixed effects and a separate cubic age polynomial for each education group, restricted to be flat at age 40.383838These are the same set of worker characteristics as in Card et al. (2013). In line with our clustering strategy, we estimate wage premiums at the group level using the group identifiers obtained from the k-means algorithm.
As shown in Table 1, a minority of employers satisfy the stability and mobility requirements in our sample. Across samples, these conditions leave us with around 25 percent of the initial number of employers. But because of the long tail in the employer-size distribution, our estimation sample still retains 80 percent of the initial number of workers and worker-year observations.
@@ 4.4 Results
Figure 4 shows two-way scatter plots of the parameter estimates in the baseline model. Starting with the joint distribution of the estimated overall employer values $V_{j}$ and employer pay premiums $\psi_{j}$ in Panel A, the key object of interest in Sorkin (2018), we find that the correlation between these model components is 0.36.393939Sorkin (2018) finds a correlation of 0.53 between overall values $V_{j}$ and pay premiums $\psi_{j}$ based on US matched employer-employee data (see Table II in his paper). Besides a different institutional setting and time period, US data on job-to-job transitions is available only at quarterly frequency and lacks information on hours worked, while we use the exact start and end date of each spell and hourly wages. Next, Panel B shows the estimated overall employer values $V_{j}$ along the horizontal axis and the corresponding employer non-pay values $a_{j}$ along the vertical axis. We also find a positive association between $V_{j}$ and employer non-pay values $a_{j}$, especially at the lower end of the value distribution. By comparison, differences in pay premiums are the only source of employer heterogeneity in the canonical wage-posting model Burdett and Mortensen (1998), which implies $a_{j}=0$ at all employers, representing a useful benchmark to compare our results. In Panel C, we find a strong positive correlation of 0.68 between $V_{j}$ and the flow utility $u_{j}$ associated with being employed at $j$, capturing both pay and non-pay values, as shown in equation (4). In Panel D, we show the joint distribution of $V_{j}$ and job security value $s_{j}=1-\delta_{j}-\rho_{j}$, finding a strong positive correlation of 0.58. By contrast, in Burdett and Mortensen (1998), this correlation is also zero by construction since job security is assumed to be the same across all employers.
Overall, panels A-D show that there is a positive gradient between the value of employers $V_{j}$ and its pay, non-pay, and job security components. The plots also suggest that these relationships are to some degree non-linear. Differences in non-pay and job security are mostly a characteristic of employers at the lower end of the $V_{j}$ distribution, while differences in pay are mostly a characteristic of employers at the upper end of the $V_{j}$ distribution.
@@
Note: The figure provides two-way scatter plots depicting our baseline model estimates for employers’ overall value ($V_{j}$), flow utility value ($u_{j}$), pay value ($\psi_{j}$), non-pay value ($a_{j}$), and job security value ($s_{j}$). The size of the scatter plot markers are proportional to the number of worker-years in each cluster. Local linear regressions and correlations are weighted by worker-years. Dashed horizontal lines refers to the standard Burdett and Mortensen (1998) model, i.e., no non-pay value and common job security across employers. The latter is calculated as the (weighted) average of job security ($s_{j}$) across employers.
Next, we focus on the joint distributions between pay premiums $\psi_{j}$ and the other sources of employer heterogeneity that determine overall employer values in our model. Importantly, in Panel E, we find a strong negative correlation of $-0.64$ between the estimated employer pay $\psi_{j}$ and non-pay $a_{j}$ values, reflecting sizable compensating differentials in our framework. By comparison, in the Burdett and Mortensen (1998) model, there are no systematic non-pay differences across employers ($\text{Corr}(\psi_{j},a_{j})=0$). Since the correlation is clearly less than zero, there is still substantial variation in non-pay $a_{j}$ at every estimated level of employer pay $\psi_{j}$.
Further, in Panel G, we show that the estimated job security component $s_{j}$ and pay component $\psi_{j}$ are positively related, with a correlation of 0.3.404040This finding is in line with the estimates in Jarosch (2023), who estimates a positive correlation between match productivity and job security in a related random search framework with job-to-job transitions. In the canonical model, this correlation is also zero by assumption since job security is the same at all employers. Finally, we find that employer flow utilities $u_{j}$ are positively correlated with pay and non-pay values in Panels F and H, respectively, consistent with the findings in Panels A-C, while there does not appear to be a clear association between employer job security $s_{j}$ and employer non-pay and flow utility as shown in Panels I and J.
@@
Overall Value Job Security Value Non-Pay Value
$V_{j}$ $s_{j}$ $a_{j}$
(1) (2) (3)
Panel A. Baseline Model: $G=\text{round}(J/50)$
   Pooled sample 0.356 0.297 -0.639
   Women 0.193 0.307 -0.712
   Men 0.469 0.256 -0.579
   College 0.317 0.301 -0.634
   Non-college 0.333 0.299 -0.654
Panel B. Smaller Clusters: $G=\text{round}(J/25)$
   Pooled sample 0.334 0.255 -0.641
   Women 0.183 0.276 -0.708
   Men 0.430 0.204 -0.590
   College 0.296 0.252 -0.615
   Non-college 0.303 0.235 -0.630
Panel C. Larger Clusters: $G=\text{round}(J/100)$
   Pooled sample 0.371 0.323 -0.635
   Women 0.190 0.327 -0.721
   Men 0.500 0.299 -0.560
   College 0.343 0.358 -0.636
   Non-college 0.346 0.302 -0.645
Notes: The table shows key summary statistics of the estimated model parameters, namely the correlation between employer pay value and overall value (Corr($\psi_{j}$,$V_{j}$)), job security value (Corr($\psi_{j}$,$s_{j}$)), and non-pay value (Corr($\psi_{j}$,$a_{j}$)), respectively, across alternative specifications and sample restrictions. The model estimates are for the period 2015-2019.

### v2

@@ 4.1 Model
Voluntary worker flows from $j$ to $k$ are given by the measure of workers employed at $j$ ($L_{j}$), who are neither forced to reallocate (with probability $\rho_{j}$) nor forced to move to non-employment (with probability $\delta_{j}$), search for an alternative employer (with probability $\lambda_{1}$), and get an offer from employer $k$ with probability $f_{k}$.252525 The probabilities $\rho_{j}$ and $\delta_{j}$ are assumed to be mutually exclusive. In Equation (2), workers choose whether to remain with employer $j$ or move to $k$, which occurs with probability $\text{Pr}(k\succ j)$. This is the sense in which these flows are voluntary and reveal workers’ preferences over the set of employers $\mathcal{J}$. We follow Sorkin (2018) and assume there are i.i.d. taste shocks $\{\varepsilon_{i}\}_{i\in\{\mathcal{J},N\}}$ that make the value of working at employer $j$ and of non-employment worker-specific in each period. We further assume that these shocks are drawn from a Gumbel distribution with location parameter normalized to zero and scale parameter $\sigma^{-1}$. [^26] Normalizing the location parameter to zero is without loss of generality because it shifts the value of working at each employer $j$ and of being in non-employment by the same amount. With this assumption, the acceptance probability for a worker with a job offer from employer $k$ currently working at $j$ is given by
@@
We depart from Sorkin (2018) by making an explicit functional form assumption on the utility flow $u_{j}$ of working at employer $j$ [^27] Sorkin (2018) makes a similar assumption on the value $V_{j}$ of employers; Lagos (2024) also follows a similar approach. By contrast, $V_{j}$ is not additively separable into pay and non-pay components in our formulation. We assume that $u_{j}$ depends linearly on a pay component $\ln W_{j}$ and a non-pay (or amenity) component $a_{j}$
@@
The additive formulation $\ln W_{j}+a_{j}$ is standard in the literature that estimates static models of worker preferences over employers (Card et al., 2018; Lamadon et al., 2022). With an additional assumption to pin down the scale of the i.i.d taste shocks $\{\varepsilon_{i}\}_{i\in\{\mathcal{J},N\}}$, we can recover the non-pay component $a_{j}$ associated with working at employer $j$ by inverting the value function in (3), conditional on the wage $W_{j}$ paid by employer $j$. We discuss this assumption below.
@@
We complete the description of the model by noting that there are similar expressions for workers’ flows to and from non-employment (counterparts to Equation (1) and Equation (2)) and for the value of worker in non-employment (a counterpart to Equation (3)), where workers in non-employment get a draw from the job offer distribution $\{f_{j}\}$ at rate $\lambda_{0}$. These expressions follow naturally and are relegated to Appendix C.
@@ 4.2 Identification
The employer-level parameters $\{\sigma V_{j},L_{j},f_{j},\rho_{j},\delta_{j}\}_{j\in\mathcal{J}}$ and the transition parameters $\lambda_{0}$ and $\lambda_{1}$ are identified from (i) the realized mobility flows between employers $M_{jk}$, and (ii) which employers are shrinking or growing (see Sorkin, 2018, Section V.B). The notation $\sigma V_{j}$ emphasizes that worker flows pin down the value of working at employer $j$ up to the scale of the idiosyncratic taste shocks, which is retrieved separately in the next steps.
We briefly outline the main source of identification for each parameter intuitively and refer to Sorkin (2018) for details. The main source of identification for $\sigma V_{j}$ is the aggregation of the realized mobility flows $M_{jk}$. An estimate for $L_{j}$ is obtained directly from the data as the share of worker-year observations at employer $j$. $f_{j}$ is identified from the share of non-employment flows to employer $j$. The separation shocks $\delta_{j}$ and $\rho_{j}$ are identified from the separation flows at shrinking firms, which give information on involuntary worker flows. Finally, the transition parameters $\lambda_{0}$ and $\lambda_{1}$ are pinned down by accounting for the aggregate transition flows, respectively from non-employment to employment and between employers.
Two sample restrictions are required for identification at this step. First, the aggregation of mobility flows only identifies $V_{j}$ within the strongly connected sets of employers.282828 Within a strongly connected set of employers, each employer has at least one worker moving in and one worker moving out. This condition is required for the fixed-point associated with the appropriately scaled matrix of worker flows to exist. See, e.g., Jackson (2010) for details on social network definitions. Second, each employer must hire at least one worker from non-employment for its sampling weight $f_{j}$ to be well-defined.
@@
where $X_{it}$ are exogenous worker characteristics (such as age), $\alpha_{i}$ is the unobserved worker effect, $j(i,t)$ is worker $i$’s employer in period $t$, and $e_{it}$ is an idiosyncratic shock. $\psi_{j(i,t)}$ is only identified within the connected sets of employers, which is already required to identify $V_{j}$ from worker flows (as a strongly connected set is connected by definition). Specifically, we assume that employers post piece-rate wage contracts as part of the utility bundle $V_{j}$. Piece-rate contracts are a commonly used assumption in random search models with wage-posting. [^30] See, for instance, Barlevy (2008) and Engbom and Moser (2022). This assumption implies that the wage premium paid by employers is log-additive in workers’ experience ($X_{it}$), ability ($\alpha_{i}$), and a time-varying and non-persistent productivity shock ($e_{it}$). Equation (6) is then correctly specified as long as individual-specific productivity does not interact with mobility decisions. We therefore also require that (i) individual-specific productivity is log-linear in the flow value of non-employment, and (ii) time-varying productivity shocks $e_{it}$ are mean-independent conditional on the realization of i.i.d. taste shocks. These assumptions ensure that individual-specific productivity does not affect the choice to move between employers and between employment and non-employment, in line with the model of mobility described so far.
@@
Conditional on $u_{j}$ and a measure for $\ln W_{j}$, the non-pay parameters can be inferred from Equation (4). To recover $u_{j}$, however, we also need to pin down the parameter $\sigma$ that controls the scale of the i.i.d. taste shocks. We do this by assuming the moment condition $\text{Var}(u_{j})=\text{Var}\big{(}\ln W_{j}\big{)}$, so the variance of flow utility is the same as the variance in the flow utility of pay $\ln W_{j}$. An intuitive justification for this moment condition is that the flow utility of pay can be expected to be an important component of the flow utility of a job, so their variance should be of similar magnitude. However, this moment condition is not without loss of generality because it restricts the correlation between $\ln W_{j}$ and $a_{j}$ to be non-positive. It is immediate to check that $\text{Var}(u_{j})=\text{Var}\big{(}\ln W_{j}\big{)}$ implies $\text{Corr}(\ln W_{j},a_{j})\leq 0$. In this specific sense, the non-pay parameters we recover can be interpreted as having a compensating differentials component by definition.
@@ 4.3 Estimation
Our sample restrictions follow from the model structure and the conditions required for identification. In our baseline estimation, we restrict the sample to the period 2015-2019 and retain workers aged 20 to 60 (inclusive). In additional analyses, we estimate the model separately for each of the remaining three five-year periods, using the same sample restrictions as in our baseline. We also restrict attention to relatively larger and more stable employers. Specifically, we require each establishment to be observed in at least two separate years and to have at least five non-singleton workers on average in each year, where a non-singleton worker is a worker observed at some later point within the sample period. In addition, the identification conditions outlined in Section 4.2 impose extra restrictions on the set of establishments included in the estimation sample. We restrict the sample to the largest strongly connected set of establishments and impose that each of these establishments hire at least one individual from non-employment. [^32] Focusing on the largest connected set is common in the literature. See, among others, Card et al. (2013) and Bonhomme et al. (2020) in the context of the two-sided unobserved heterogeneity regression model (6) and Sorkin (2018) and Morchio and Moser (2024) in the context of the fixed-point revealed preferences model. These last two restrictions are interdependent, and they are imposed recursively until the set of establishments converges.
@@
We obtain estimates of the establishment wage premium from the two-sided unobserved heterogeneity regression (6). We estimate this equation separately in each sample. Our wage measure is log average earnings per hour, where earnings are total pre-tax earnings and hours are total contract hours over the duration of each yearly employment spell. The vector of worker-level characteristics $X_{it}^{\prime}$ includes year-by-education fixed effects and a separate cubic age polynomial for each education group, restricted to be flat at age 40.353535These are the same set of worker characteristics as in Card et al. (2013). In line with our clustering strategy, we estimate wage premiums at the cluster level using the group identifiers obtained from the k-means algorithm.
As shown in Table 1, a minority of employers satisfy the stability and mobility requirements. Across samples, these conditions leave us with around 25 percent of the initial number of establishments. However, due to the long tail of the employer-size distribution, our estimation sample still retains around 75 percent of the initial number of workers.
@@ 4.4 Results
Figure 4 shows two-way scatter plots of the parameter estimates in the baseline model. Starting with the joint distribution of the estimated overall employer values $V_{j}$ and employer pay premiums $\psi_{j}$ in Panel A, the key object of interest in Sorkin (2018), we find that the correlation between these model components is 0.36.363636Sorkin (2018) finds a correlation of 0.53 between overall values $V_{j}$ and pay premiums $\psi_{j}$ based on US matched employer-employee data (see Table II in his paper). Besides a different institutional setting and time period, US data on job-to-job transitions is available only at quarterly frequency and lacks information on hours worked, while we use the exact start and end date of each spell and hourly wages. Next, Panel B shows the estimated overall employer values $V_{j}$ along the horizontal axis and the corresponding employer non-pay values $a_{j}$ along the vertical axis. We also find a positive association between $V_{j}$ and employer non-pay values $a_{j}$, especially at the lower end of the value distribution. By comparison, differences in pay premiums are the only source of employer heterogeneity in the canonical wage-posting model Burdett and Mortensen (1998), which implies $a_{j}=0$ at all employers, representing a useful benchmark to compare our results. Taken together, the preceding results indicate a strong positive correlation between $V_{j}$ and the flow utility $u_{j}$, which captures both pay and non-pay values, as shown in Equation (4). In Panel C, we show the joint distribution of $V_{j}$ and job security value $s_{j}=1-\delta_{j}-\rho_{j}$, finding a strong positive correlation of 0.55. By contrast, in Burdett and Mortensen (1998), this correlation is also zero by construction since job security is assumed to be the same across all employers.
Overall, panels A-C show that there is a positive gradient between the value of employers $V_{j}$ and its pay, non-pay, and job security components. The plots also suggest that these relationships are to some degree non-linear. Differences in non-pay and job security are mostly a characteristic of employers at the lower end of the $V_{j}$ distribution, while differences in pay are mostly a characteristic of employers at the upper end of the $V_{j}$ distribution.
@@
Note: The figure provides two-way scatter plots depicting our baseline model estimates for employers’ overall value ($V_{j}$), flow utility value ($u_{j}$), pay value ($\psi_{j}$), non-pay value ($a_{j}$), and job security value ($s_{j}$). The size of the scatter plot markers are proportional to the number of worker-years in each cluster. Local linear regressions and correlations are weighted by worker-years. Dashed horizontal lines refers to the standard Burdett and Mortensen (1998) model, i.e., identical job security across employers and no amenities. We calculated the former as the weighted average of $s_{j}$ across employers.
Next, we focus on the joint distributions between pay premiums $\psi_{j}$ and the other sources of employer heterogeneity that determine overall employer values in our model. Importantly, in Panel D, we find a strong negative correlation of $-0.58$ between the estimated employer pay $\psi_{j}$ and non-pay $a_{j}$ values, reflecting sizable compensating differentials in our framework. By comparison, in the Burdett and Mortensen (1998) model, there are no systematic non-pay differences across employers ($\text{Corr}(\psi_{j},a_{j})=0$). Since the correlation is clearly less than zero, there is still substantial variation in non-pay $a_{j}$ at every estimated level of employer pay $\psi_{j}$.
Further, in Panel F, we show that the estimated job security component $s_{j}$ and pay component $\psi_{j}$ are positively related, with a correlation of 0.29.373737This finding is in line with the estimates in Jarosch (2023), who estimates a positive correlation between match productivity and job security in a related random search framework with job-to-job transitions. In the canonical model, this correlation is also zero by assumption since job security is the same at all employers. Finally, we find that employer flow utilities $u_{j}$ are positively correlated with pay and non-pay values in Panels E and G, respectively, consistent with the findings in Panels A-C, while there does not appear to be a clear association between employer job security $s_{j}$ and employer non-pay as shown in Panel H.
@@
Overall Value Job Security Value Non-Pay Value Number of Clusters
$V_{j}$ $s_{j}$ $a_{j}$ $G$
(1) (2) (3) (4)
Panel A: Baseline Model: $G=\text{round}(J/50)$
   Pooled Sample 0.357 0.294 -0.582 1,563
   Women 0.200 0.309 -0.653 850
   Men 0.474 0.262 -0.518 858
   College 0.314 0.307 -0.591 762
   Non-College 0.339 0.291 -0.590 937
Panel B: Smaller Clusters: $G=\text{round}(J/25)$
   Pooled Sample 0.339 0.258 -0.589 3,125
   Women 0.185 0.266 -0.652 1,699
   Men 0.443 0.247 -0.538 1,717
   College 0.294 0.255 -0.597 1,525
   Non-College 0.314 0.230 -0.595 1,873
Panel C: Larger Clusters: $G=\text{round}(J/100)$
   Pooled Sample 0.366 0.326 -0.577 781
   Women 0.192 0.343 -0.659 425
   Men 0.497 0.275 -0.503 429
   College 0.333 0.311 -0.580 381
   Non-College 0.343 0.317 -0.589 468
Notes: The table shows key summary statistics of the estimated model parameters, namely the correlation between employer pay value and overall value (Corr($\psi_{j}$,$V_{j}$)), job security value (Corr($\psi_{j}$,$s_{j}$)), and non-pay value (Corr($\psi_{j}$,$a_{j}$)), respectively, across alternative specifications and sample restrictions. The model estimates are for the period 2015-2019. We weight by the number of worker-year observations in each cluster.